if the difference of two numbers is 10 and their product is 18,what is the sum of their squares?
Answer by checkley71(8403)

X-Y=10 X=10+Y
XY=18
(10+Y)Y=18
10Y+Y^2=18
Y^2+10Y-18=0
Y=(10+-11.557)/2
WHEN
Y=-11.557
THEN
X-(-11.557)=10
X=10-11.577
X=-1.557
AND
WHEN
Y=1.557
THEN
X-1.577=10
X=10=1.557
X=11.577